# eMoE-Tracker: Environmental MoE-based Transformer for Robust Event-guided Object Tracking

### v1

@@ IV-A Experimental Settings
Our method is trained end-to-end on 1 NVIDIA A800 GPU with PyTorch implementation. During training, our method utilizes a global batch size of 64 and takes 60 epochs each epoch processes $6\times 10^{4}$ sample pairs. We employ the AdamW [42] optimizer with a weight decay of $10^{-4}$ and set the initial learning rate as $2\times 10^{-4}$ but decreasing after 32 epochs by the factor of 10.
@@ IV-B Comparison
TABLE III: Ablation studies on the effectiveness of our proposed modules: eMoE and CRM.
Model eMoE CRM VisEvent COESOT
@@ IV-C Ablation Studies
Inserted layers of eMoE We achieve RBG-E tracking under various challenging conditions by injecting visual prompt blocks into different layers of the backbone model. It is intuitive to investigate the effect on the number of prompt blocks. Here we set different insert intervals for blocks to insert and the intervals are 1,2 4, 6, and 12. Therefore, the first means that all the layers are fully inserted and the last one only inserts the blocks in the last high-level layer.
@@
TABLE VI: Model complexity Comparison between two stream trackers with our eMoE-Tracker. OS and TS denote one stream and two streams, respectively.
Tracker AFNet [7] FENet [7] VisEvent [8] eMoE-Tracker
Structure-type TS TS TS OS
Trainable Parameters(MB) 25.16 41.87 27.53 8.42
@@ VIII-B Attributes Performance
Since the VisEvent and COESOT dataset all provide 17 attributes for tracking performance evaluation, we can leverage the Matlab toolkit to plot the diagram of curves about the attributes performance comparison. We represent the diagram of curves on 16 attributes except for the no motion scenario in Fig. 10 and Fig. 9, which show the superior performance under all the environmental conditions compared to other SOTA priors.

### v2

@@ IV-A Experimental Settings
Our method is trained end-to-end on 1 NVIDIA A800 GPU with PyTorch implementation. During training, our method utilizes a global batch size of 64 and takes 60 epochs, and each epoch processes $6\times 10^{4}$ sample pairs. We employ the AdamW [42] optimizer with a weight decay of $10^{-4}$ and set the initial learning rate as $2\times 10^{-4}$ but decreasing after 32 epochs by the factor of 10.
@@ IV-B Comparison
TABLE III: Ablation studies on the effectiveness of our proposed modules: eMoE and CRM, and tracking header frozen or not.
Model eMoE CRM Header Unfrozen VisEvent COESOT
@@ IV-C Ablation Studies
Inserted layers of eMoE. We achieve RBG-E tracking under various challenging conditions by injecting eMoE blocks into different layers of the backbone model. It is intuitive to investigate the effect on the number of inserted blocks. Here we set different insert intervals for blocks to insert and the intervals are 1,2 4, 6, and 12. Therefore, the first means that all the layers are fully inserted and the last one only inserts the blocks in the last high-level layer.
@@
TABLE VI: Trainable parameters comparison among two stream trackers, one stream trackers, and our eMoE-Tracker. OS and TS denote one stream and two streams, respectively.
Tracker AFNet [7] FENet [7] VisEvent [8] ViPT [10] CEUTrack [4] eMoE-Tracker
Structure-type TS TS TS OS OS OS
Trainable Parameters(MB) 25.16 41.87 27.53 0.84 93.7 8.42
@@ VIII-A Visualization
TABLE VII: Testing Result in FELT-SOT dataset.
Dataset SR PR NPR
FELT-SOT [wang2024FELTSOT] 65.1 72.9 73.6
@@ VIII-B Attributes Performance
Attention Map. Since the VisEvent and COESOT dataset all provide 17 attributes for tracking performance evaluation, we can leverage the Matlab toolkit to plot the diagram of curves about the attributes performance comparison. We represent the diagram of curves on 16 attributes except for the no motion scenario in Fig. 11 and Fig. 10, which show the superior performance under all the environmental conditions compared to other SOTA priors.
Feature Map. To validate that the disentangled environmental experts can learn discriminative attribute-specific features and assemble them in proper weights, we illustrate features and weights from four environmental experts in Fig. 9.
Figure 9: The visualization of the feature maps and assembling weights for four environmental experts branches under specific scenarios. The 4-digits vector is the attributes label and the following numbers are the corresponding learnable weights for attribute-specific features. The last column denotes the assembled features.
@@
VIII-C Real-World Testing
For real-world testing, we further validate our method on another RGB-event dataset, namely VOT-DVS, to demonstrate the practicality. The VOT-DVS dataset is a popular synthesized RGB-event tracking dataset. The testing result is reported in Table. VII.